OpenStudy (anonymous):

Marco needed to boil some water in order to cook his pasta. One minute after he began heating the water, it reached the temperature of 95 degrees. Ten minutes after boiling the water, the temperature rose to 212 degrees. Write an equation to describe its meaning in standard form.

OpenStudy (anonymous):

Do you know how to find the slope of that line?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

 Here's a rough sketch of what the line looks like.

OpenStudy (anonymous):

Slope is 'rise-over-run' or vertical change divided by horizontal change. Find that first.

OpenStudy (anonymous):

would that be 11.7?

OpenStudy (anonymous):

No, but that's close. What numbers did you divide to get that?

OpenStudy (anonymous):

why is it linear?

OpenStudy (anonymous):

We only have 2 points, @satellite73 , no choice but to do linear interpolation.

OpenStudy (anonymous):

is the slope 13?

OpenStudy (anonymous):

It's a lousy assumption, but not much else to go on here.

OpenStudy (anonymous):

Yes, the slope is 13. Good job! Now you can use one of those points, and the slope to get an equation. You can use slope-intercept form, y=mx+b, or point-slope form, y-y1=m(x-x1) to make the equation, then rearrange it to get standard form.

OpenStudy (anonymous):

in slope- intercepy form, would the equation be y=13x+82?

OpenStudy (anonymous):

intercept*

OpenStudy (anonymous):

That seems to check out. You can check it by putting in your x,y points to see if it works.

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

You're welcome. The question asks for standard form, which for a line is Ax+By=C, where A, B, and C are integers.